How can I solve the equation
(y-11)- (y+6) = 8y

Answer by stanbon(75887)

solve the equation
(y-11)- (y+6) = 8y
y-11-y-6 = 8y
-17 = 8y
y = -17/8
===============
Cheers,
Stan H.